### v2

# ReGATE: Learning Faster and Better with Fewer Tokens in MLLMs

Abstract
The computational cost of training multimodal large language models (MLLMs) grows rapidly with the number of processed tokens. Existing efficiency methods mainly target inference via token reduction or merging, offering limited benefits during training. We introduce ReGATE (Reference-Guided Adaptive Token Elision), an adaptive token pruning method for accelerating MLLM training. ReGATE adopts a teacher-student framework, in which a frozen teacher LLM provides per-token guidance losses that are fused with an exponential moving average of the student’s difficulty estimates. This adaptive scoring mechanism dynamically selects informative tokens while skipping redundant ones in the forward pass, substantially reducing computation without altering the model architecture. Across three representative MLLMs, ReGATE matches the peak accuracy of standard training on MVBench up to 2 faster, using only 38% of the tokens. With extended training, it even surpasses the baseline across multiple multimodal benchmarks, cutting total token usage by over 41%. Code and models will be released publicly.
1 Introduction
Multimodal large language models (MLLMs) face significant challenges due to the high computational cost of training [jin2024efficient]. A key bottleneck is the self-attention mechanism, whose complexity grows quadratically with input sequence length [nips2017_attention]. This issue is especially severe in video tasks [videosavi, videopasta, avatar], where frames are tokenized into extremely long sequences. Consequently, training MLLMs on large-scale datasets requires substantial computing resources, limiting accessibility and slowing progress. Improving training efficiency is therefore essential for enabling MLLMs to scale to longer contexts, support richer modalities, and learn from larger datasets that would otherwise be infeasible to process.
Several strategies have been proposed to speed up inference in MLLMs, including token pruning [ye2024atpllava, arif2025hired, alvar2025divpruned, lee2025tamp], token merging [chen2024fastv, dhouib2025pact], and token compression [yang2025libra]. While these methods reduce FLOPs during inference, they rely on pre-scored or frozen tokens and cannot adapt token importance dynamically during training. As a result, reducing training costs remains a more complex and less explored problem. In unimodal text models, learnable token pruning methods such as RHO-1 [lin2024rho1] have improved training efficiency, but these approaches have not been extended to large multimodal models. Early efforts to speed up visual processing, often focusing on standard vision transformers [akbari2021vatt] or early video-language models [lei2021clipbert], rely on heuristics like random token dropping. In multimodal settings, however, determining which tokens are important often depends on subtle visual evidence, temporal cues, or interactions between multiple modalities. Ignoring these dependencies can lead to the removal of crucial information, resulting in inefficient training and reduced multimodal understanding. This gap highlights the need for a principled method to estimate token importance during training, one that captures cross-modal dependencies without relying on heuristics.
To address this challenge, we introduce ReGATE (Reference‑Guided Adaptive Token Elision), a framework designed to accelerate the training of MLLMs (Figure 1). ReGATE adopts a teacher-student architecture, in which the student is the multimodal model being trained, and the teacher is a frozen, text-only version of the same LLM backbone. ReGATE combines two complementary signals to identify and retain the most informative tokens during training dynamically. First, it checks whether a token requires visual grounding by seeing if the text-only teacher can predict it from the prompt alone. Second, it evaluates the student model’s learning progress using an exponential moving average (EMA) of token-wise historical losses. By integrating these signals, ReGATE allocates computation to the subset of tokens that are both critical for multimodal understanding and remain challenging for the model to learn. To validate the effectiveness and generality of ReGATE, we evaluate it on three representative MLLMs: VideoChat2 [li2024mvbench], VideoLLaMA2 [cheng2024videollama2], and InternVL3.5 [wang2025internvl35], covering a wide range of architectures, training paradigms, and scales, and assess performance across diverse image and video benchmarks.
To summarize, our contributions are threefold:
We introduce ReGATE, an adaptive token pruning method for accelerating MLLM training. ReGATE leverages a text-only reference teacher model and the student’s historical token difficulty to dynamically identify and retain visually essential tokens, without introducing any additional trainable parameters.
We show that the model-agnostic ReGATE integrates seamlessly into existing MLLMs, requiring no architectural changes, making it easy to adopt.
Extensive experiments on five image benchmarks and eight video benchmarks demonstrate ReGATE’s broad applicability and efficiency. On the challenging MVBench benchmark, ReGATE reaches the baseline’s peak accuracy up to 2 faster while using only 38% of the tokens on average.
2 Related Work
2.1 Token Compression for Fast Inference
Most existing work focuses on accelerating inference, not training. Inference-time sparsity methods show that many tokens can be removed or merged with minimal impact on accuracy [shi2025staticdynamic, jiang2025visa, yang2025topv, hyun2025multigranular]. In vision transformers, Dynamic Token Pruning [tang2023dtop] halts processing of easy tokens layer by layer, reducing FLOPs by 20–35% without degrading performance. For video LLMs, DyCoke [tao2025dycoke] compresses spatial-temporal tokens during inference, achieving up to 2 speed-ups while keeping model weights frozen. Moving beyond pruning, Importance-Based Token Merging [wu2025imagebased] merges similar tokens rather than dropping them, maintaining performance on long-video benchmarks while delivering 1.5 faster inference. However, all these methods operate after training is complete. During training, the full token is still processed in every forward and backward pass, leaving the computational cost of training mainly unaddressed.
2.2 Token Compression for Fast Training
Only a few studies have explored token compression during training, rather than just at inference. In text‑only language models, RHO‑1 [lin2024rho1] ranks tokens with a reference model and backpropagates only through the most difficult subset, reducing pre‑training tokens by 50% while improving accuracy. For MLLMs, LaVi [yue2025lavi] injects vision-conditioned deltas into layer norms to skip visual tokens but needs extra modulation layers. LLaVA‑Meteor [li2025llavameteror] introduces a flash‑fusion module and a dual‑expert scorer that prunes 75–95% visual tokens during instruction tuning but adds extra parameters and targets only visual tokens. In contrast, ReGATE combines a static cross-modal reference loss from a text-only teacher with a dynamic EMA-based student signal. Together, these signals produce an adaptive, parameter-free sparsity mechanism that selectively gates textual tokens while preserving visual ones, without modifying the model architecture.
2.3 Teacher‑Student Distillation for MLLMs
Most distillation approaches for MLLMs mainly focus on parameter compression [li2023unlock, cai2025llavakd, udandarao2025activedatacuratione]. A systematic study [xu2024llavadi] shows that jointly aligning tokens and logits helps a smaller student model inherit visual grounding from a larger teacher model. Similarly, methods like DIME‑FM [sun2023dimefm] show how cross-modal features can be transferred even from unpaired data. More recently, MaskedKD [son2024rolemasking] reduces teacher FLOPs by up to 50% by masking patch tokens based on student attention, but it still backpropagates through all student tokens. In contrast, ReGATE leverages the teacher’s per-token loss to decide which tokens the student should process during each forward and backward pass. This approach significantly reduces computation without altering parameters and enables model-agnostic sparsity for more efficient training.
3 ReGATE
We introduce ReGATE, a method that speeds up the training of MLLMs by selectively allocating computational resources only to tokens that truly require visual information. The key insight is that not all tokens in a multimodal sequence depend equally on visual context: some can be accurately predicted from text alone, while others need cross-modal grounding. To capture this, ReGATE uses a teacher-student framework. The student is the main MLLM being trained. The teacher is a reference model created by taking the student’s LLM backbone, removing its visual components (the visual encoder and projector), and freezing its weights. This results in a pure text-only LLM that acts as a fixed expert to estimate the degree to which each token depends on visual input. Given a batch of input sequences containing both text and visual tokens, we generate a binary mask that determines which tokens should be actively computed and which can be skipped. This section explains how we calculate per-token difficulty scores using the frozen text-only teacher combined with the student’s own training history, how we dynamically adjust the fraction of tokens retained during training, and how we apply the resulting mask within the transformer decoder.
3.1 Difficulty Score Formulation
Let denote the token sequence in sample , including both text tokens and special visual tokens (e.g., <image> or <video> tokens representing visual content). To compute the reference loss, we construct a modified sequence by replacing the actual visual tokens with placeholder tokens (typically the padding token <pad>), ensuring the sequence length remains identical to the original multimodal input fed to the MLLM’s backbone LLM. Our reference model is a pure text-only LLM obtained by removing the visual encoder and projector from the MLLM backbone, thus incapable of processing any visual content. By feeding the constructed placeholder sequence to the reference model in evaluation mode, we compute the per-token negative log-likelihood:
A low value of indicates that the teacher can predict based on the textual context alone, whereas a high value signals that multimodal information is needed to predict the token. In parallel, we monitor how difficult each token has been for the student across training updates. For every training sample and token position , we maintain a running difficulty buffer updated as an exponential moving average (EMA) of the student’s cross-entropy loss:
where is the current cross-entropy loss of the student model at token position , and controls the smoothing of the EMA. A higher value of indicates that token in sample has consistently posed difficulties during training. We then combine the reference loss and the student’s historical difficulty into a unified score for each token:
where balances these two signals. Tokens with a higher combined difficulty, , either consistently challenge the student model or genuinely require visual context, and thus are prioritized during the training updates. Note that this combined difficulty evaluation is performed exclusively on output tokens (labels), as these tokens directly influence the training process through backpropagation.
3.2 Dual‑cycle Sparsity Schedule
We employ a deterministic schedule to determine the fraction of tokens kept at each training step. Our schedule repeats every steps. In the first steps of each cycle, we keep all tokens (i.e., ) to allow the model to stabilize. In the remaining steps, we retain only a fixed proportion of the tokens. Formally, if denotes the global training step, we have:
3.3 Dynamic Token Gating
For each sample , we identify the indices of valid tokens excluding padding and special markers. Let denote those indices and . We compute the combined difficulty for using Equation (3) and select the top tokens. The resulting binary mask is set to one for retained tokens and zero otherwise. We always retain all special visual tokens (e.g., those corresponding to a frame or image) regardless of their difficulty to preserve multimodal information. Because the difficulty buffer is updated after every epoch, the set of selected positions adapts throughout training: tokens that become easy for the student are gradually deprioritised, while persistently challenging tokens or those requiring visual grounding remain active. This dynamic gating enables the model to allocate its computational budget to the most informative parts of the sequence at each epoch, rather than committing to a fixed sparsity pattern. Finally, the per‑sample binary masks are concatenated and padded to form a batch mask where is the expanded sequence length accounting for visual tokens.
3.4 Adaptive Decoder Sparsity
To exploit the binary mask during forward propagation, we modify the transformer decoder layer of the backbone LLM. We implement sparse attention by passing the mask directly as the attention mask to flash attention routines and by zeroing out the hidden states of pruned tokens. For the feed‑forward network, we gather only the active positions, apply the MLP to them, and scatter the outputs back to their positions. The residual connections ensure that skipped tokens retain their previous representations. A single decoder layer forward pass applies self-attention and MLP only to tokens selected by the binary mask, while skipped tokens are propagated unchanged via residual connections. This design introduces no additional parameters, integrates seamlessly with libraries such as HuggingFace Transformers, and remains fully compatible with pre-trained weights. During back propagation, gradients are computed only for parameters associated with the kept tokens. Skipped tokens are treated as constants. Their activations and losses receive no gradient updates, while residual connections ensure that gradient flow remains continuous across layers. This masking rule aligns the backward graph with the sparse forward path and maintains training stability in practice. Complete pseudocode and implementation details (including Q/K/V masking) are provided in Supplementary Material.
4 Experiments
| Model | LLM | Tokens | ScienceQA | MME | VizWiz | POPE | |
[…]
| InstructBLIP [dai2023instructblip] | Vicuna-7B | – | 60.5 | 254.3/1137.1 | 34.5 | 86.1 | 46.4 |
| LLaVA-1.5 [liu2024llava] | Vicuna-7B | – | 66.8 | 302.1/1506.2 | 50.0 | 85.9 | 66.1 |
| Qwen-VL-Chat [bai2023qwenvl] | Qwen-7B | – | 68.2 | 392.1/1467.8 | 38.9 | 74.9 | 58.2 |
| LLaVA-1.6 [liu2024llava] | Vicuna-7B | – | 70.1 | – | 57.6 | 86.5 | 70.2 |
| VILA1.5 [lin2024vila] | Llama-2-13B | – | 79.1 | 288.9/1429.3 | 60.6 | 84.2 | 62.8 |
| LLaVA-Next [liu2024llavanext] | Mistral-7B | – | 73.0 | 308.9/1512.3 | – | 87.3 | 72.4 |
| LLaVA-OneVision [llavaov] | Qwen2-7B | – | 95.4 | 415.7/1577.8 | 53.0 | 87.4 | 75.4 |
| Qwen2.5-VL [bai2025qwen25vl] | Qwen2.5-7B | – | 89.0 | 613.9/1698.1 | – | 85.9 | 77.0 |
[…]
| Claude3.7-Sonnet [anthropic2025claude37] | – | – | 90.9 | 649.6/1189.7 | – | 82.4 | 74.3 |
| Gemini-1.5-Flash [geminiteam2024gemini15] | – | – | 83.3 | 488.6/1589.3 | – | 88.5 | 75.0 |
| Gemini-1.5-Pro [geminiteam2024gemini15] | – | – | 85.7 | 548.2/1562.4 | – | 88.2 | 76.0 |
| GPT-4o [openai2024gpt4ocard] | – | – | 90.1 | 719.3/1609.4 | – | 85.0 | 76.4 |
| GPT-4.1 [openai2024gpt4ocard] | – | – | 92.8 | 673.9/1663.6 | – | 86.4 | 78.0 |
[…]
| VideoChat2 | Mistral-7B | 3.93B | 40.8 | 314.6/1244.0 | 28.5 | 86.2 | 45.9 |
[…]
| VideoLLaMA2 | Qwen2-7B | 83.82M | 61.4 | 376.4/1474.0 | 46.8 | 86.7 | 70.4 |
[…]
| InternVL3.5 | Qwen3-14B | 3.96B | 93.3 | 681.6/1694.3 | 60.6 | 91.6 | 76.8 |
[…]
4.1 Implementation Details
To demonstrate the effectiveness and generality of ReGATE, we apply it to three representative MLLMs, VideoChat2 [li2024mvbench], VideoLLaMA2 [cheng2024videollama2], and InternVL3.5 [wang2025internvl35], spanning diverse architectures, training paradigms, and scales. We exclude models such as Qwen2.5-VL [bai2025qwen25vl] and VideoLLaMA3 [zhang2025videollama3] due to the lack of publicly available pretrained weights. Training such models from scratch is impractical, as they rely on web-scale data and hundreds of GPUs. However, with sufficient resources and pretrained weights, ReGATE can be seamlessly integrated into any MLLM training pipeline. For all experiments, the per-token reference losses from the text-only teacher are computed once over the entire fine-tuning dataset before student training begins and then cached and reused throughout training.
VideoLLaMA2. We first apply ReGATE to VideoLLaMA2-7B [cheng2024videollama2], whose Qwen2-7B [yang2024qwen2] backbone is unfrozen during multimodal fine-tuning. ReGATE is introduced at this stage, using a text-only reference teacher (the same backbone without the visual encoder or adapter) that computes per-token losses with visual tokens masked out. This configuration allows us to examine how ReGATE interacts with full fine-tuning on a standard multimodal backbone.
VideoChat2. To evaluate ReGATE under parameter-efficient fine-tuning (PEFT), we integrate it into the LoRA-based Stage 3 training of VideoChat2-7B [li2024mvbench], built on Mistral-7B [jiang2023mistral7b]. Gradients for LoRA parameters are computed only from the high-importance tokens selected by ReGATE, while the language backbone weights remain frozen. This setup enables us to assess whether ReGATE can complement PEFT methods without architectural modifications.
InternVL3.5. Finally, we evaluate scalability of ReGATE on InternVL3.5-14B [wang2025internvl35], based on Qwen3-14B [yang2025qwen3technicalreport]. The reference teacher is derived from the same backbone by removing its visual encoder and projector, forming a text-only LLM. During fine-tuning, ReGATE dynamically gates tokens to focus computation on the most informative positions, reducing activation memory and training FLOPs. This setup allows us to assess ReGATE’s scalability and stability at large model scales.
Datasets and sparsity schedule. We fine-tune VideoLLaMA2 with and without ReGATE on the VideoChatGPT dataset [maaz2024videochatgpt], which is a subset of VideoLLaMA2’s official fine-tuning dataset containing approximately 300,000 instruction-response pairs. For VideoChat2, we similarly use a subset of its official fine-tuning data comprising around 2.6 million instruction pairs. For InternVL3.5, we adopt the same 2.6M-sample dataset used for VideoChat2, since the full InternVL3.5 training corpus is extremely large and not fully released. Training follows the dual-cycle sparsity schedule described in Section 3.2, with parameters set to , , and . To ensure stable training at the start, we prepend a global warm-up phase of 100 iterations, during which all tokens are retained. The main hyperparameters for ReGATE include an exponential moving average (EMA) decay of and a teacher loss weighting coefficient of . All experiments on VideoLLaMA2 and VideoChat2 are run on 4 H100 GPUs, while experiments on InternVL3.5 use 16 H100 GPUs, all under mixed-precision training.
4.2 Benchmarks and Baselines
We evaluate ReGATE across image, long-video, and short-video understanding benchmarks under the LMMs-Eval [zhang2024lmmseval] framework. For image understanding, we include ScienceQA [lu2022scienceqa], MME [fu2024mme], VizWiz [gurari2018vizwiz], POPE [li2023pope], and SEED [li2024seed], which together probe scientific reasoning, perception, answerability, hallucination, and broad multimodal comprehension. For long-video understanding, we evaluate on Video-MME [fu2025videomme], LongVideoBench [wu2024longvideobench], MLVU [zhou2025mlvu], and EgoSchema [mangalam2023egoschema], all of which emphasize extended temporal reasoning, consistency, and grounding across minutes- to hour-long sequences. For short-video understanding, we adopt MVBench [li2024mvbench], Perception Test [pătrăucean2023perceptiontest], Vinoground [zhang2024vinoground], and NExT-QA [xiao2021nextqa], which target fine-grained event recognition, local temporal relations, counterfactual reasoning, and causal/temporal action reasoning in clips of tens of seconds.
We evaluate ReGATE against a diverse set of state-of-the-art models, including high-performing open-source families (LLaVA, Qwen) and proprietary models (Google Gemini, OpenAI GPT, Anthropic Claude). This selection spans multiple LLM backbones and sizes, ensuring robust comparisons across image and video tasks (Tables 4, 4.2, LABEL:tab:short-video).
| Model | LLM | Frames | Tokens | LongVideoBench | MLVU | EgoSchema | |
[…]
| Video-LLaVA [lin2024videollava] | Vicuna-7B | 8 | – | 39.9 | 39.1 | 47.3 | 38.4 |
| LLaMA-VID [li2024llamavid] | Llama-2-7B | 1fps | – | 25.9 | – | 33.2 | 38.5 |
| LLaVA-NeXT-Video [zhang2024llavanextvideo] | Vicuna-7B | 32 | – | – | 43.5 | – | 43.9 |
| LLaVA-NeXT-Video [zhang2024llavanextvideo] | Qwen2-32B | 32 | – | 60.2 | – | 65.5 | 60.9 |
| VILA1.5 [lin2024vila] | Llama-2-40B | 8 | – | 60.1 | – | 56.7 | 58.0 |
| LLaVA-OneVision [llavaov] | Qwen2-7B | 32 | – | 58.2 | 56.4 | 64.7 | 60.1 |
| Qwen2.5-VL [bai2025qwen25vl] | Qwen2.5-7B | – | – | 65.1 | 56.0 | 70.2 | 65.0 |
| VideoLLaMA3 [zhang2025videollama3] | Qwen2.5-7B | 1fps | – | 66.2 | 59.8 | 73.0 | 63.3 |
[…]
| Gemini-1.5-Flash [geminiteam2024gemini15] | – | – | – | 70.3 | 61.6 | – | 65.7 |
| Gemini-1.5-Pro [geminiteam2024gemini15] | – | – | – | 75.0 | 64.0 | – | 71.2 |
| GPT-4o [openai2024gpt4ocard] | – | – | – | 71.9 | 66.7 | 64.6 | 72.2 |
[…]
| VideoChat2 | Mistral-7B | 16 | 3.93B | 26.0 | 21.8 | 36.0 | 55.6 |
[…]
| VideoLLaMA2 | Qwen2-7B | 16 | 83.82M | 53.7 | 47.7 | 53.2 | 58.2 |
[…]
| InternVL3.5 | Qwen3-14B | 16 | 3.96B | 62.4 | 57.9 | 63.7 | 64.7 |
[…]
4.3 Results
Learning better: ReGATE’s accuracy gains across image and video benchmarks. The comprehensive results presented in Tables 4, 4.2, and LABEL:tab:short-video show how VideoLLaMA2, VideoChat2, and InternVL3.5 perform, with and without ReGATE, across a range of image, short video, and long video understanding benchmarks. ReGATE improves performance consistently by focusing computation on the most informative tokens. Figure LABEL:fig:saliency visualizes the learned attention maps, showing that ReGATE concentrates attention on visually and semantically important tokens compared to standard fine-tuning. For example, VideoLLaMA2-ReGATE outperforms the baseline VideoLLaMA2 on most tasks while using 41.22% fewer tokens. Similarly, VideoChat2-ReGATE achieves better results than the baseline VideoChat2 while using 43.51% fewer tokens. At the larger 14B scale, InternVL3.5-ReGATE also shows gains across both image and video benchmarks, confirming that ReGATE remains effective for high-capacity backbones.
On image understanding tasks requiring multimodal reasoning, all three models show significant gains. VideoLLaMA2-ReGATE improves by 19.1% on ScienceQA and by up to 33.1 points on MME. VideoChat2-ReGATE improves by 5.8% and 46.1 points on the same benchmarks. InternVL3.5-ReGATE also shows clear improvements, including +7.7/+4.5 points on MME, 0.9% on VizWiz, and 1.6% on POPE. For long-video understanding, VideoChat2-ReGATE shows strong improvements of 6.7% on Video-MME and 4.5% on MLVU. VideoLLaMA2-ReGATE also improves, though more modestly, with gains of 0.8% and 1.3% on the same tasks. At the larger scale, InternVL3.5-ReGATE further improves performance by 0.6% on Video-MME, and 0.5% on MLVU. Short video tasks benefit as well. VideoLLaMA2-ReGATE improves by 1.6% on MVBench and 1.1% on Perception, and VideoChat2-ReGATE gains 0.9% and 1.6%. InternVL3.5-ReGATE also shows improvements, with +1.3% on MVBench and +1.4% on Perception. Beyond performance metrics, ReGATE generalizes well across architectures and training strategies. It works effectively with transformer backbones such as Mistral, Qwen2, and Qwen3, and retains its benefits under both full-parameter and LoRA fine-tuning. These results confirm that ReGATE’s efficiency and accuracy gains are architecture- and training-strategy–agnostic, enabling seamless integration into diverse multimodal models.
Supplementary Material
6 Sparse Decoder Forward and Masking Details
Algorithm 1 shows how the binary mask controls computation within each decoder layer. The hidden states are first normalized, and self-attention is applied only to tokens marked active in the mask. The outputs of attention are added back through the residual connection, after which the MLP is executed solely on the active positions. Finally, the updated values are written back to the corresponding slots, while masked tokens remain unchanged and are propagated forward through residual paths. This ensures that only informative tokens are processed, while inactive ones are efficiently skipped without disrupting model integrity.
Q/K/V Masking. For clarity, we specify how the binary mask is applied within transformer layers. During self-attention, queries (Q) are computed only for the active tokens, while keys (K) and values (V) are restricted to the same set, with pruned tokens excluded from the gather step. In the feed-forward network (FFN), hidden states at active positions are passed through the MLP and the outputs are written back to their original slots. Residual connections ensure that skipped tokens simply propagate unchanged across layers. This design realizes sparse computation while leaving the model parameters intact and fully compatible with standard transformer implementations.
7 Additional Ablation Studies
In this section, we delve deeper into the design choices of ReGATE, specifically examining the sensitivity of the weighting hyperparameter and the impact of the reference teacher’s capacity.
7.1 Impact of the Weighting Factor
To validate the contributions of the individual components within our dual-signal token scoring mechanism, we conduct an ablation study on the hyperparameter . This coefficient balances the two core signals in our difficulty score formulation: , where is the student’s dynamic EMA difficulty and is the static reference loss from the teacher model. We evaluate three values for : 0.0, 0.5, and 1.0. The experiments use our VideoLLaMA2-ReGATE setup with all other hyperparameters fixed for a fair comparison.
As shown in Table 7, setting achieves the best balance. Relying solely on the student’s historical difficulty () fails to capture zero-shot visual dependencies, while relying exclusively on the teacher () ignores the student’s training dynamics. The combined signal effectively isolates tokens that are both visually critical and challenging for the current model state.
| Description | Acc. (%) | |
| Student EMA Only | 47.7 | |
| Reference Loss Only | 46.4 | |
| Combined Signals | 48.9 |
7.2 Impact of Reference Teacher Capacity
| Teacher Model | Avg. Acc. (%) |
| Qwen2-1.5B | 45.4 |
| Qwen2-57B | 46.8 |
| Qwen2-7B (Ours) | 48.9 |
A critical design choice in ReGATE is the selection of the reference teacher. We investigate how the capacity of the teacher model affects the student’s performance. Specifically, for the VideoLLaMA2-7B [cheng2024videollama2] student (based on Qwen2-7B [yang2024qwen2]), we evaluate reference signals from a smaller model (Qwen2-1.5B) and a substantially larger model (Qwen2-57B-A14B), comparing them against the matched Qwen2-7B baseline. To ensure fair comparison, we restrict the study to a single model family, which guarantees consistent tokenization and avoids misalignment between teacher and student. Cross-architecture teachers (e.g., Mistral guiding Qwen) are excluded because distinct tokenizers yield inconsistent token boundaries for the same text, violating the one-to-one mapping required for per-token reference loss computation in ReGATE. Such mismatch renders the reference signal mathematically invalid.
The results in Table 8 show that the matched Qwen2-7B teacher outperforms both the smaller 1.5B and the much larger 57B models. This pattern reveals a capacity mismatch effect: performance degrades when the teacher’s capability diverges too far from the student’s.
When using the large Qwen2-57B teacher, performance drops because the teacher is too strong relative to the student. With its vast world knowledge, the 57B model can often predict complex tokens purely from textual context, producing very low reference loss. However, the 7B student lacks this prior knowledge and instead depends on visual grounding to infer these tokens. As a result, the large teacher mistakenly signals ReGATE to prune these positions, causing the student to skip crucial computations. Conversely, the smaller Qwen2-1.5B teacher underestimates the student’s ability. It assigns high loss to linguistic patterns or factual details that the 7B student already handles easily. This floods the top- selection with linguistically difficult but visually irrelevant tokens, wasting computation on trivial cases instead of focusing on visual reasoning.
Overall, these findings suggest that the optimal reference teacher should be capacity-aligned with the student. The reference loss is most effective when it reflects the student’s own uncertainty, directing computation precisely where textual priors fall short.
| Benchmark | Response formatting prompts |
| POPE | – |
| MME | Answer the question using a single word or phrase. |
| VisWiz | Answer the question using a single word or phrase. When the provided information is insufficient, respond with “Unanswerable”. |
| ScienceQA | Answer with the option’s letter from the given choices directly. |
| SEED | Answer with the option’s letter from the given choices directly. |
| MLVU | – |
| MVBench | Only give the best option. |
| Video-MME | Answer with the option’s letter from the given choices directly. |
| EgoSchema | Answer with the option’s letter from the given choices directly. |
| NExT-QA | – |
| Perception | Answer with the option’s letter from the given choices directly. |
| Vinoground | Please only output one English character. |
| LongVideoBench | Answer with the option’s letter from the given choices directly. |
8 Additional Benchmarks Details
Table 9 lists the evaluation prompts corresponding to each benchmark used in the experiments, most of which are adapted from LMMs-Eval [zhang2024lmmseval].
9 Qualitative Analysis of Reference Loss
To validate the core mechanism of ReGATE, we qualitatively analyze the reference loss signal that guides its token selection. We assume that a high loss score from the text-only teacher indicates that a token requires visual information to be understood. Figure 5 shows two video Q&A examples, visualizing the loss for each word in the answer as calculated by a Mistral-7B [jiang2023mistral7b] teacher model.
The results strongly support our assumption. As illustrated in the figure, tokens for visual details that are hard to guess from text alone, like the action “mixing” or the attribute “reflective”, get high loss scores. In contrast, simple grammatical words like “The” and “is”, or terms repeated from the question like “bartender”, get low scores. This difference confirms that reference loss is a reliable indicator of visual importance, enabling ReGATE to focus its computation on the most critical tokens for more efficient training.